Predation. Answer questions 1&2

Faced with a variety of potential food choices, pr
Answer #1

1.E/(Th+TS)=12.26kj/3.1s+37.9=0.299 for Chitons

E/(Th+TS)=1.42kj/2.9s+18.9=0.0651 for Mussels

       As the energy value of chiton is higher than Mussel even after reducing the value by half so it should opt for Chiton.

2.Natural Selection will always be beneficial for the predator when the prey is source of high energy but the availibility also matters and high availibity of the prey will only be possible when it is fit to survive in the enviornment.

by evolution only the survival f fittest matters.